OpenStudy (anonymous):

Find the area between y=4x and y=x^2?

OpenStudy (anonymous):

here the graph y= 4x cuts the graph y=x^2 at 2 points. we can find the points by equating the 2 graph. 4x = x^2 x=0 & x= 4 let the area under the 1st graph is a & area under the 2nd graph is b. then the required area is a-b. u can find the areas by integration .

OpenStudy (anonymous):

hey u have to find the area from x=0 to x=4 only in both case. then u can do the other manipulation .

OpenStudy (anonymous):

so what do you get for the area?

OpenStudy (anonymous):

by integration. \[a=\int\limits_{0}^{4}4x dx =2x ^{2} =2*4*4-0 =32\] \[b=\int\limits_{0}^{4} x ^{2}dx = x^{3}/3 = 4^{3}/3-0 =64/3\] area = a-b = (32-64/3) = 32/3